11 10. JUNE di Toy - ruze , YU -- m 11. Find the Laplace transform inverse of e-** 2015
Answer #1

Solution F18) = .-38 2s+5 e 82_48+13 =e-58 208-21 +5+2x2 (8-2)2+(3)2 = 2(8-2) e Is 3. 3e-58 08-2)2+(372 (8-2)2+13)2 take inve